The bright red color in fireworks is due to emission of light when Sr(NO3)2 is heated. If the wavelength is 650.0 nm, what is the frequency of this light?
Answer #1

For a given wavelength of light:

Frequency of the light= Velocity of light / wavelength of the light.

Velocity of light= 3 ×108m/s

Given wavelength = 6.5 × 10-7m

Frequency = 3×108 / 6.5 × 10-7 s-1

= 4.6 × 1014 s-1